OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3 ln x - 2 ln c

zepdrix (zepdrix):

Here is an important rule of logarithms:\[\huge a \log(b) \qquad = \qquad \log(b^a)\] How will this rule help us rewrite \(\large 3 \ln x\) ?

OpenStudy (anonymous):

is the answer in x^3/c^2

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

okay thanks